A satellite orbits the earth a distance of 1.50 × 107 m above the planet's surface and takes 8.65 hours for each revolution about the earth. The earth's radius is 6.38 × 106 m. The acceleration of this satellite is closest to


1
Expert's answer
2021-05-25T16:16:16-0400

The acceleration in a circular motion is defined as:

"a = \frac{v^{2}}{r} \;(1)"

Where a is the centripetal acceleration, v the velocity and r is the radius.

The equation of the orbital velocity is defined as

"v = \frac{2 \pi r}{T}\; (2)"

Where r is the radius and T is the period

For this particular case, the radius will be the sum of the high of the satellite"(1.50 \times 10^{7} \;m)" and the Earth radius "(6.38 \times 10^{6} \;m)" :

"r = 1.50 \times 10^{7} \;m+6.38 \times 10^{6} \;m \\

r = 21.38 \times 10^{6} \;m"

Then, equation 2 can be used:

"T = 8.65 \;hrs \times \frac{3600 \;s}{1\;hrs}=31140 \; s \\

v = \frac{2 \pi (21.38 \times 10^{6}\;m)}{31140 \;s} \\

v = 4313 \;m/s"

Finally equation 1 can be used:

"a = \frac{(4313 \;m/s)^{2}}{21.38 \times 10^{6}m} \\

a = 0.87\; m/s^{2}"

Hence, the acceleration of the satellite is "0.87 \;m/s^{2}"